Belcanto Corporation experienced a fire on December 31, 2017, in which its financial records were partially destroyed. It has been able to salvage some of the records and has ascertained the following balances. Cash Accounts receivable (net) Inventory Accounts payable Notes payable Common stock, $100 par Retained earnings December 31, 2017 40,000 84,000 200,000 50,000 30,000 400,000 170,000 December 31, 2016 15,000 126,000 180,000 10,000 20,000 400,000 101,000 Additional information 1. The inventory turnover is 4.2 times. 2. The return on common stockholders equity is 14%. The company had no additional paid-in-capital 3. The accounts receivable turnover is 10.2 times 4. The return on assets is 12.5% 5. Total assets, Dec. 31, 2016 604,750Additional information: 1. The inventory turnover is 4.2 times 2. The return on common stockholders equity is 14%. The company had no additional paid-in-capital. 3. The accounts receivable turnover is 10.2 times 4. The return on assets is 12.5%. 5. Total assets, Dec. 31, 2016 604,750 Compute the following values for 2017 (a) Cost of goods sold s (b) Net credit sales$ (c) Net income (d) Total assets
Answer #1

a) Cost of goods sold :

Average inventory = (200000+180000/2) = 190000

Cost of goods sold = 190000*4.2 = $798000

b) Net credit sales

Average receivable = (84000+126000/2) = 105000

Net credit sales = 105000*10.2 = $1071000

c) Net income :

Average common Stockholder's equity = (400000+400000/2) = $400000

Net income = 400000*14% = $56000

d) Total assets

Average total assets = 56000/12.5 = $448000

Total assets 2017 = (448000*2)-604750 = $291250